OpenStudy (anonymous):

Finding the equation of the line that is tangent to the following curve at (1, pi/2): it then gives the equation: 2xy+pi sin y=2pi What rules do i use to find the derivative?

OpenStudy (anonymous):

Use the chain rule.

OpenStudy (anonymous):

\[ 2xy+\pi \sin y=2\pi \]First do \[ (2xy)' \]This requires product rule.

OpenStudy (anonymous):

Next do \(\pi\sin y\), this just requires chain rule.

OpenStudy (anonymous):

\[ 2xy+\pi \sin y=2\pi \\ (2xy)'+(\pi \sin y)' = 0 \]

OpenStudy (anonymous):

Need help?

OpenStudy (anonymous):

sorry! was working on helping someone. one second :D

OpenStudy (anonymous):

the deriative of 2xy is 2. but what do i do with 2pi?

OpenStudy (anonymous):

Nope \[ (2xy)' = (2x)' \cdot y + 2x\cdot y' = 2y+2xy' \]

OpenStudy (anonymous):

whaaat. ok so, the deriative of (2x)'=2 why does y change things?

OpenStudy (anonymous):

product rule!

OpenStudy (anonymous):

Okay suppose \(f(x) = 2x\) and \(g(x) = y\)

OpenStudy (anonymous):

I'm just saying \[ (fg)' = f'g+fg' \]

OpenStudy (anonymous):

duh. im stupid.

OpenStudy (anonymous):

Now can you do \[ (\pi \sin y)' \]

OpenStudy (anonymous):

yay thank you. It might take me a few minutes. Id be grateful if you checked back :D

OpenStudy (anonymous):

I'm not going to check your work unless you show me the steps though.

OpenStudy (anonymous):

Of course. \[\frac{ d }{ dy }(\pi \sin (y))= \pi(\frac{ d }{dy })\sin(y)= \pi(\frac{ d }{ dy })\cos(y)= \pi \cos(y)\]

OpenStudy (anonymous):

But you want to differentiate with respect to \(x\)

OpenStudy (anonymous):

I have no idea how to do that.

OpenStudy (anonymous):

use chain rule: \[ \frac{d}{dx}f(y) = f'(y)y'(x) \]

OpenStudy (anonymous):

Basically\[ \frac{d}{dx}f(y) = \frac{df}{dy}\frac{dy}{dx} = f'(y)y' \]

OpenStudy (anonymous):

You found \(f'(y)\) already.

OpenStudy (anonymous):

\[ f'(y) = \pi \cos y \]

OpenStudy (anonymous):

oh ok i thought you were asking me to use the chain rule to find the derivative of pi sin y lol

OpenStudy (anonymous):

was trying to wrap my head around that

OpenStudy (anonymous):

\[ \frac{d}{dx}\pi \sin y = (\pi \cos y)y' \]

OpenStudy (anonymous):

so \[\pi \cos(y)\] is correct then?

OpenStudy (anonymous):

Notice the \(y'\)

OpenStudy (anonymous):

ok. i see it. why is it there? :(

OpenStudy (anonymous):

Okay.... Look chain rule says: \[ [f(g(x)) ]' = f'(g(x)) g'(x) \]

OpenStudy (anonymous):

Let \(y = g(x)\) and you get: \[ \frac{d}{dx}f(y) = f'(y) y' \] In this case \( f(y) = \pi \sin y \)

OpenStudy (anonymous):

is the \[f'= 2y+2xy'?\]

OpenStudy (anonymous):

... I'm tired dude...

OpenStudy (anonymous):

lol. sorry. its fine. thank you for your help.

OpenStudy (anonymous):

I can show you my answer...

OpenStudy (anonymous):

no thats okay. I'd rather figure it out first.

OpenStudy (anonymous):

I dont understand how to apply the chain rule to different equations, im doing practice problems to attempt to understand it better, But i appreciate the effort!

OpenStudy (anonymous):

\[ 2xy+\pi \sin y=2\pi \\ (2xy)'+(\pi \sin y)' = 0 \\ 2y+2xy'+(\pi \cos y)y' = 0 \\ (2x+\pi \cos y)y' = -2y \\ y' = -\frac{2y}{2x+\pi \cos y} \]